OpenStudy (anonymous):

pleaseee help, how can I simplify (- 2q^4/7)^4 ?

OpenStudy (anonymous):

\[\frac{-2q^{16}}{7^4}\] Now solve..

OpenStudy (anonymous):

That wouldn't be the answer?

OpenStudy (anonymous):

Remember this rule.. \[(a^m)^n=a^{mn}\]

OpenStudy (anonymous):

I'm so confused man.

OpenStudy (anonymous):

??

OpenStudy (anonymous):

Have you got any answers ?.

OpenStudy (anonymous):

No... I'm really confused

OpenStudy (anonymous):

According to the rule.. \[{(-2q^4)^4}\]

OpenStudy (anonymous):

2q^16

OpenStudy (anonymous):

-2q^16*

OpenStudy (anonymous):

\[=-2q^{16}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Yeahhhh..

OpenStudy (anonymous):

No what?

OpenStudy (anonymous):

Now*

OpenStudy (anonymous):

The denominator is also like that..

OpenStudy (anonymous):

ok so -2q^16/7^4 but how do I solve that?

jhonyy9 (jhonyy9):

so i think first of all you need deciding 7 is denominator of exponent or ????

OpenStudy (anonymous):

What do you mean?

jhonyy9 (jhonyy9):

-2q^4 (------ )^4 7 so hence will be right or how ????

OpenStudy (anonymous):

It's not even an answer choice in my question. -2q^4/7 is

OpenStudy (anonymous):

I mean -2q^16/7

jhonyy9 (jhonyy9):

so than sign that 7 will be the indice of radical yes ?

OpenStudy (anonymous):

I have no idea what you're talking about.

jhonyy9 (jhonyy9):

so this is because x^(1/2) =sqrt(x) so 2 is indice of radical ok ?

jhonyy9 (jhonyy9):

@Miyuru do you understand it ?